At present, the number of government personnel in NCR is estimated at 450,000. If it gains 2.3% of this number by accession and loses 1.2% of this number by separation regularly each year for 5 years. How many government personnel are there in NCR at the end of the period? ( answer is 474,750 show your solution )
Answer by richwmiller(17219)

The wording is odd. Why would you keep using 450000 as the basis each year?
450000+5*(.023*450000)-5*(.012*450,000)=
450000+51750-27000=
450000+24750=
474750